# A Study on the Calibration of In-context Learning

### v3

@@ Abstract
Accurate uncertainty quantification is crucial for the safe deployment of language models (LMs), and prior research has demonstrated improvements in the calibration of modern LMs. Our study focuses on in-context learning (ICL), a prevalent method for adapting static LMs through tailored prompts, and examines the balance between performance and calibration across a broad spectrum of natural language understanding and reasoning tasks. Through comprehensive experiments, we observe that, with an increasing number of ICL examples, models initially exhibit increased miscalibration before achieving better calibration and miscalibration tends to arise in low-shot settings. Moreover, we find that methods aimed at improving usability, such as fine-tuning and chain-of-thought (CoT) prompting, can lead to miscalibration and unreliable natural language explanations, suggesting that new methods may be required for scenarios where models are expected to be reliable.
@@ 1 Introduction
Language models (LMs) that encompass transformer-based architectures (Brown et al., 2020; Chowdhery et al., 2023; OpenAI, 2023) can generate coherent and contextually relevant texts for various use cases. Despite their impressive performance, these models occasionally produce erroneous or overconfident outputs, leading to concerns about their calibration (Dawid, 1982; DeGroot and Fienberg, 1983) which measures how faithful a model’s prediction uncertainty is. Such a problem is pressing when users adapt them using a recent paradigm called in-context learning (Brown et al., 2020) to construct performant predictors, especially for applications in safety-critical domains (Bhatt et al., 2021; Kadavath et al., 2022; Pan et al., 2023).
We provide an in-depth evaluation and analysis of how well these models are calibrated - that is, the alignment between the model’s confidence in its predictions and the actual correctness of those predictions. This token-level calibration assessment enables us to measure the discrepancy between the model’s perceived and actual performance through a Bayesian uncertainty lens, providing a valuable metric for assessing the model’s accuracy and reliability.
Refer to caption (a) Demonstration of In-context Learning
@@
We find that LM such as LLaMA (Touvron et al., 2023a) is poorly calibrated in performant settings and there exists a calibration-accuracy trade-off (Fig.1) for low-shot settings ($k<4$), i.e. as we increase the amount of in-context samples, both prediction accuracy and calibration error increase. Such a trade-off can be improved using more ICL examples ($k=8$) and larger models. Crucially, this calibration degradation worsens when fine-tuning occurs using specialized data to improve usability, such as curated instructions (Dubois et al., 2023), dialogues (Zheng et al., 2023), or human preference data (Ziegler et al., 2019). Though previous work (Braverman et al., 2020) shows the entropy of each generation step is drifting and can be recalibrated via scaling techniques (Platt et al., 1999) such as temperature scaling (Guo et al., 2017), we show that in contrast to classic regimes, the miscalibration issue in ICL can not be easily addressed using such well-established recalibration approaches that rely on additional validation data.
Moreover, we study the trade-off in reasoning tasks that involve generation of explanations (Camburu et al., 2018; Nye et al., 2021; Wei et al., 2022) before the answer, showing that the model can produce confidently wrong answers (using confidence histograms and reliability plots) when prompted with explanations on Strategy QA (Geva et al., 2021), Commonsense QA (Talmor et al., 2018), OpenBook QA (Mihaylov et al., 2018), World Tree (Jansen et al., 2018). We carefully design our human evaluation and observe that, with the increase in model sizes and the quantity of ICL examples, there is a corresponding rise in the proportion of confidently predicted examples among those incorrectly forecasted. Moreover, we find that a high proportion of wrong predictions are of high confidence and showcase those typical confidently wrong examples of LMs.
In-context learning has been hypothesized as meta-learning by gradient descent in their forward pass (Von Oswald et al., 2023), which might hopefully yield calibrated predictions (Błasiok et al., 2023) if the models are getting close to local optimality with respect to test loss through meta-optimization. However, the fact that choosing ICL samples from the validation set does not naturally lead to calibrated predictions shows that ICL learns in a fairly different way than SGD. We design controlled experiments to illustrate the task learning properties of ICL, showing that when examples in the prompt demonstrate consistent task properties, the learning performance would be improved.
@@ 2 Related Work
Calibration of language models. Calibration is a safety property to measure the faithfulness of machine learning models’ uncertainty, especially for error-prone tasks using LLMs. Previous works find that pre-training (Desai and Durrett, 2020) and explanation (Zhang et al., 2020; González et al., 2021) improves calibration. Models can be very poorly calibrated when we prompt LMs (Jiang et al., 2021), while calibration can also depend on model size (Kadavath et al., 2022). (Braverman et al., 2020) assesses the long-term dependencies in a language model’s generations compared to those of the underlying language and finds that entropy drifts as models such as when GPT-2 generates text. The intricacy of explanations on complementary team performance poses additional challenges due to the overreliance on explanations of users regardless of their correctness (Bansal et al., 2021). (Mielke et al., 2022) gives a framework for linguistic calibration, a concept that emphasizes the alignment of a model’s expressed confidence or doubt with the actual accuracy of its responses. The process involves annotating generations with <DK>, <LO>, <HI> for confidence levels, then training the confidence-controlled model by appending the control token <DK/LO/HI> at the start of the output, followed by training a calibrator to predict these confidence levels, and finally predicting confidence when generating new examples. (Tian et al., 2023) finds that asking LMs for their probabilities can be better than using conditional probabilities in a traditional way. LHTS (Shih et al., 2023) is a simple amortized inference trick for temperature-scaled sampling from LMs and diffusion models. To aggregate log probabilities across semantically equivalent outputs, Kuhn et al. (2023) utilize bidirectional entailment through a model to identify outputs that are semantically similar, thereby refining the uncertainty estimation process. (Cole et al., 2023) identifies the calibration challenge in ambiguous QA and distinguishes uncertainty about the answer (epistemic uncertainty) from uncertainty about the meaning of the question (denotational uncertainty), proposing sampling and self-verification methods. (Kamath et al., 2020) trains a calibrator to identify inputs on which the QA model errs and abstains when it predicts an error is likely. (Zhao et al., 2023) proposes the Pareto optimal learning assessed risk score for calibration and error correction but requires additional training. (Kalai and Vempala, 2023) show the trade-off between calibration and hallucination but they didn’t study it in a realistic setting and how the predicted answer’s accuracy would impact those two safety aspects.
@@ 3 Background
Table 1: Accuracy and Calibration of LLaMA-30B model with three sizes across four text classification datasets and four reasoning datasets. Results are excluded when the data exceeds the context length limit.
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@
Empirical estimate of the expected calibration error (ECE) In the realm of probabilistic classifiers, calibration is a crucial concept. A classifier, denoted as $\mathcal{P}_{\theta}$ with parameters $\theta$ and operating over $C$ classes, is said to be "canonically calibrated" when, for every probability distribution $p$ over the $C$ classes and for every label $y$, the probability that the label is $y$ given the classifier’s prediction is $p$ matches the component of $p$ corresponding to $y$. This is mathematically represented as:
$\forall p\in\Delta^{C-1},\forall y\in Y:P\left(Y=y\mid\mathcal{P}_{\theta}(X)=p\right)=p_{y}.$
@@
A simpler calibration criterion is the "top-label calibration." In this case, a classifier is deemed calibrated if, for every top predicted probability $p^{*}$, the probability that the true label belongs to the class with the highest predicted probability, given that this maximum predicted probability is $p^{*}$, equals $p^{*}$. Formally:
$\forall p^{*}\in[0,1]:P\left(Y\in\arg\max p\mid\max\mathcal{P}_{\theta}(X)=p^{*}\right)=p^{*}.$
To gauge the calibration of a model, we adopt Expected Calibration Error (ECE) defined as:
$\mathbb{E}\left[\left|p^{*}-\mathbb{E}\left[Y\in\arg\max\mathcal{P}_{\theta}(X)\mid\max\mathcal{P}_{\theta}(X)=p^{*}\right]\right|\right].$
@@ 4.2 Numerical Results
Performance of LLaMA. We record the performance and calibration errors for $k$-shot learning ($k=0,1,2,3,4,8$), characterizing the calibration-accuracy trade-off in both classic and realistic settings (Tab. 1). Our findings are two-fold: as more in-context examples are included, we observe a concurrent rise in both accuracy and calibration error across most low-shot situations. Especially, when $k=0$ increases to $k=1$, there is a marked boost in both accuracy and calibration error, demonstrating the importance of in-context examples in learning performance while one single example may not be able to reduce aleatoric uncertainty. Moreover,
In particular, for reasoning tasks, we explore prompting approaches that explicitly include explanations in reasoning tasks, i.e. scratchpad (Nye et al., 2021) or chain-of-thought (Wei et al., 2022), showing that calibration significantly degrades after generating a long context for reasoning and explaining the final answer.
@@
Post-hoc recalibraiton. We experiment with three strategies (Algorithm 1) in applying temperature scaling methods (Guo et al., 2017) to fix miscalibration:
@@
    (00-shot) Learn a temperature from the training split and apply it to all test samples with different shot numbers;
@@
    ($k$-shot) We learn one temperature for each $k$-shot ICL, i.e., we learn different temperatures for different shot numbers in ICL;
@@
    (Fix $w$) For each experiment, we fix the prompt and learn the temperature for the fixed prompt. That is, for every possible ICL prompt, we learn a corresponding temperature for calibration.
@@
The training algorithm is shown in Appendix Alg. 1. Looking into Tab. 3, none of the above strategies achieves satisfactory calibration performance, which is in contrast to the well-studied supervised learning setting where scaling the confidence scores can effectively reduce calibration errors (Guo et al., 2017). The fact that applying a post-processing calibration method, such as temperature scaling, cannot directly resolve the miscalibration issue suggests that ICL might have substantially different properties compared to making predictions via classical supervised learning models.
Refer to caption (a) Classification accuracy Refer to caption (b) Calibration error
@@
Figure 3: Accuracy and calibration errors of base models LLaMA and Mistral, as well as fine-tuned variants. The reported values for Acc and ECE represent averages across experiments conducted with 0, 1, …, 8 shots.
@@
Table 3: ECE for different calibration strategies of base models LLaMA-7B across various shot settings.
Dataset Strategy Shot Avg
0 1 2 3 4 8
SST-2 Vanilla 0.038 0.098 0.133 0.136 0.124 0.064 0.099
00-shot 0.026 0.157 0.265 0.275 0.282 0.222 0.205
$k$-shot 0.026 0.104 0.103 0.101 0.087 0.035 0.076
Fix $w$ 0.027 0.079 0.056 0.051 0.051 0.030 0.049
CB Vanilla 0.133 0.274 0.164 0.148 0.192 0.206 0.186
00-shot 0.126 0.272 0.143 0.142 0.183 0.177 0.174
$k$-shot 0.126 0.247 0.137 0.153 0.203 0.251 0.186
Fix $w$ 0.138 0.183 0.127 0.141 0.150 0.194 0.155
RTE Vanilla 0.130 0.103 0.122 0.147 0.130 0.083 0.119
00-shot 0.117 0.100 0.130 0.154 0.130 0.094 0.121
$k$-shot 0.114 0.106 0.130 0.161 0.121 0.104 0.123
Fix $w$ 0.114 0.078 0.097 0.104 0.079 0.079 0.092
AGNews Vanilla 0.086 0.077 0.166 0.203 0.088 0.063 0.114
00-shot 0.101 0.087 0.149 0.198 0.117 0.078 0.122
$k$-shot 0.114 0.085 0.161 0.203 0.075 0.067 0.117
Fix $w$ 0.132 0.081 0.113 0.126 0.089 0.079 0.103
Refer to caption (a) 00-shot Refer to caption (b) $1$-shot Refer to caption (c) $4$-shot Refer to caption (d) $8$-shot
@@
Figure 4: Illustration of confidence distribution: the number of samples whose confidence is greater than a threshold on Commonsense QA.
Refer to caption (a) 00-shot Refer to caption (b) $1$-shot Refer to caption (c) $4$-shot Refer to caption (d) $8$-shot
@@
The effect of fine-tuning. We show that vicuna, alpaca, and LLaMA2-Chat are all more accurate but less calibrated than their LLaMA counterpart backbones (Fig. 3), the margin is especially large for reasoning tasks and vicuna. Our finding indicates that fine-tuning might significantly degrade calibration, corroborating the evidence reported in GPT-4 (OpenAI, 2023), albeit it can improve the reasoning accuracy dramatically. Our results provide evidence that though fine-tuned on carefully curated datasets can greatly improve question-answering performance, especially for hard tasks like reasoning problems, attention may need to be paid when assessing the calibration of those models’ predictions. We also note that the pattern of zero-shot performance is fairly different for two fine-tuned models, i.e. vicuna, and alpaca. Moreover, we include results of Mistral-7B (Jiang et al., 2023), a sparse Mixture of Experts (MoEs) architecture with sliding window attention. As a base model, it shows similar performance and calibration compared with LLaMA2-7B, indicating that our conclusion still holds for the model pre-trained with significantly different data and architectures.
@@
The findings, as detailed in Tab. 2, reveal certain insights: firstly, integrating labels within prompts significantly decreases uncertainty and enhances learning performance. In contrast, simply repeating the context without labels does not lead to better outcomes. Secondly, the diversity of ICL examples in the prompt greatly affects performance, a potential explanation is it promotes better task learning (Pan, 2023). Those observations corroborate that ICL is performant when the number of ICL examples is large and they demonstrate consistent task properties. Importantly, the trade-off persists for different controlled scenarios, i.e. as we increase the number of ICL examples, models tend to be first more miscalibrated and then calibrated.
Table 4: Norm of representation, entropy, and confidence of LLaMA-30B model across three text classification datasets.

### v4

@@ Abstract
Accurate uncertainty quantification is crucial for the safe deployment of machine learning models, and prior research has demonstrated improvements in the calibration of modern language models (LMs). We study in-context learning (ICL), a prevalent method for adapting static LMs through tailored prompts, and examine the balance between performance and calibration across a broad spectrum of natural language understanding and reasoning tasks. Through comprehensive experiments, we observe that, with an increasing number of ICL examples, models initially exhibit increased miscalibration before achieving better calibration and miscalibration tends to arise in low-shot settings. Moreover, we find that methods aimed at improving usability, such as fine-tuning and chain-of-thought (CoT) prompting, can lead to miscalibration and unreliable natural language explanations. Furthermore, we explore recalibration techniques and find that a scaling-binning calibrator can reduce calibration errors consistently.
@@ 1 Introduction
Language models (LMs) that encompass transformer-based architectures (Brown et al., 2020; Chowdhery et al., 2023; OpenAI, 2023) can generate coherent and contextually relevant texts for various use cases. Despite their impressive performance, these models occasionally produce erroneous or overconfident outputs, leading to concerns about their calibration (Dawid, 1982; DeGroot and Fienberg, 1983) which measures how faithful a model’s prediction uncertainty is. Such a problem is pressing when users adapt them using a recent paradigm called in-context learning (Brown et al., 2020) to construct performant predictors, especially for applications in safety-critical domains (Bhatt et al., 2021; Pan et al., 2023).
We provide an in-depth evaluation and analysis of how well these models are calibrated - that is, the alignment between the model’s confidence in its predictions and the actual correctness of those predictions. This token-level calibration assessment enables us to measure the discrepancy between the model’s perceived and actual performance to assess its accuracy and reliability through a Bayesian uncertainty lens.
(a) Demonstration of In-context Learning
Refer to captionRefer to caption
@@
We find that LM such as LLaMA (Touvron et al., 2023a) is poorly calibrated in performant settings and there exists a calibration-accuracy trade-off (Fig.1) for low-shot settings ($k<4$): as we increase the amount of in-context samples, both prediction accuracy and calibration error increase. Such a trade-off can be improved using more ICL examples ($k=8$) and larger models. Crucially, this calibration degradation worsens when fine-tuning occurs using specialized data to improve usability, such as curated instructions (Dubois et al., 2023), dialogues (Zheng et al., 2023), or human preference data (Ziegler et al., 2019). Though previous common practice suggests recalibrating models’ logits via temperature scaling (Guo et al., 2017), we show that in contrast to classic regimes, the miscalibration issue in ICL can not be easily addressed using such well-established scaling approaches (Platt et al., 1999). Thus we propose to use scaling-binning (Kumar et al., 2019), which fits a scaling function, bins its outputs, and then outputs the average of the function values in that bin, to reduce the expected calibration error below $0.1$.
Furthermore, we study the trade-off in reasoning tasks that involve generation of explanations (Camburu et al., 2018; Nye et al., 2021; Wei et al., 2022) before the answer, showing that the model can produce confidently wrong answers (using confidence histograms and reliability plots) when prompted with explanations on Strategy QA (Geva et al., 2021), Commonsense QA (Talmor et al., 2018), OpenBook QA (Mihaylov et al., 2018), World Tree (Jansen et al., 2018). We carefully design our human evaluation and observe that, with the increase in model sizes and the quantity of ICL examples, there is a corresponding rise in the proportion of confidently predicted examples among those incorrectly forecasted. Moreover, we find that a high proportion of wrong predictions are of high confidence and showcase those typical confidently wrong examples of LMs.
Moreover, we find that choosing ICL samples from the validation set does not naturally lead to calibrated predictions, showing that ICL learns in a fairly different way than stochastic gradient descent, a common prototype previous works hypothesize (Von Oswald et al., 2023). Motivated by this difficulty, we design controlled experiments to illustrate that when examples in the prompt are sampled from the same task instead of repeating a given example in various ways, the learning performance would be improved.
@@ 2 Related Work
Calibration of language models. Calibration is a safety property to measure the faithfulness of machine learning models’ uncertainty, especially for error-prone tasks using LMs. Previous works find that pre-training (Desai and Durrett, 2020) and explanation (Zhang et al., 2020; González et al., 2021) improves calibration. Models can be very poorly calibrated when we prompt LMs (Jiang et al., 2021), while calibration can also depend on model size (Kadavath et al., 2022). (Braverman et al., 2020) assesses the long-term dependencies in a language model’s generations compared to those of the underlying language and finds that entropy drifts as models such as when GPT-2 generates text. The intricacy of explanations on complementary team performance poses additional challenges due to the overreliance on explanations of users regardless of their correctness (Bansal et al., 2021). (Mielke et al., 2022) gives a framework for linguistic calibration, a concept that emphasizes the alignment of a model’s expressed confidence or doubt with the actual accuracy of its responses. The process involves annotating generations with <DK>, <LO>, <HI> for confidence levels, then training the confidence-controlled model by appending the control token <DK/LO/HI> at the start of the output, followed by training a calibrator to predict these confidence levels, and finally predicting confidence when generating new examples. (Tian et al., 2023) finds that asking LMs for their probabilities can be better than using conditional probabilities in a traditional way. LHTS (Shih et al., 2023) is a simple amortized inference trick for temperature-scaled sampling from LMs and diffusion models. To aggregate log probabilities across semantically equivalent outputs, Kuhn et al. (2023) utilize bidirectional entailment through a model to identify outputs that are semantically similar, thereby refining the uncertainty estimation process. (Cole et al., 2023) identifies the calibration challenge in ambiguous QA and distinguishes uncertainty about the answer (epistemic uncertainty) from uncertainty about the meaning of the question (denotational uncertainty), proposing sampling and self-verification methods. Kamath et al. (2020) trains a calibrator to identify inputs on which the QA model errs and abstains when it predicts an error is likely. Zhao et al. (2023) proposes the Pareto optimal learning assessed risk score for calibration and error correction but requires additional training. Kalai and Vempala (2023) show the trade-off between calibration and hallucination but they didn’t study it in a realistic setting and how the predicted answer’s accuracy would impact those two safety aspects.
@@ 3 Background
Table 1: Accuracy and Calibration of LLaMA-30B model with across four text classification datasets and four reasoning datasets. Results are excluded when the data exceeds the context length limit.
@@
Empirical estimate of the expected calibration error (ECE) In the realm of probabilistic classifiers, calibration is a crucial concept. A classifier, denoted as $\mathcal{P}_{\theta}$ with parameters $\theta$ and operating over $C$ classes, is said to be "canonically calibrated" (Kull and Flach, 2015) when, for every probability distribution $p$ over the $C$ classes and for every label $y$, the probability that the label is $y$ given the classifier’s prediction is $p$ matches the component of $p$ corresponding to $y$. This is mathematically represented as, $\forall p\in\Delta^{C-1},\forall y\in Y:$
$P\left(Y=y\mid\mathcal{P}_{\theta}(X)=p\right)=p_{y}.$
@@
A simpler calibration criterion is the "confidence calibration." In this case, a classifier is deemed calibrated if, for every top predicted probability $p^{*}$, the probability that the true label belongs to the class with the highest predicted probability, given that this maximum predicted probability is $p^{*}$, equals $p^{*}$. Formally: $\forall p^{*}\in[0,1],$
$P\left(Y=c(X)\mid\max\mathcal{P}_{\theta}(X)=p^{*}\right)=p^{*},$
where $c(X)=\arg\max p$ and ties are broken arbitrarily. To gauge the calibration of a model, we adopt Expected Calibration Error (ECE Guo et al. (2017)) defined as:
$\mathbb{E}\left[\left|p^{*}-\mathbb{E}\left[Y=c(X)\mid\max\mathcal{P}_{\theta}(X)=p^{*}\right]\right|\right].$
@@ 4.2 Numerical Results
Model performance and calibration. We record the performance and calibration errors for $k$-shot learning ($k=0,1,2,3,4,8$), characterizing the calibration-accuracy trade-off in both classic and realistic settings (Tab. 1). Our findings are two-fold: as more in-context examples are included, we observe a concurrent rise in both accuracy and calibration error across most low-shot situations. Especially, when $k=0$ increases to $k=1$, there is a marked boost in both accuracy and calibration error, demonstrating the importance of in-context examples in learning performance while one single example may not be able to reduce aleatoric uncertainty. In particular, for reasoning tasks, we explore prompting approaches that explicitly include explanations in reasoning tasks, i.e. scratchpad (Nye et al., 2021) or chain-of-thought (Wei et al., 2022), showing that calibration significantly degrades after generating a long context for reasoning and explaining the final answer. We also note that having more ICL examples does not necessarily lead to better calibration though the predictive performance can generally improve (e.g., $k=8$ for CB in Tab.1). This may stem from the intrinsic limitations of transformers in effectively modeling long-term dependencies.
@@
Post-hoc recalibraiton. We conducted experiments with three strategies (Algorithm 1) to address miscalibration using temperature scaling (Guo et al., 2017) and scaling-binning (Kumar et al., 2019) with learnable parameter $w$:
@@
    (00-shot) Learning $w$ from the training split and applying it to all test samples with different shot numbers.
@@
    ($k$-shot) Learning $w$ for each $k$-shot ICL; in other words, different temperatures are learned for different shot numbers in ICL.
@@
    (Fix $w$) Fixing the prompt for each experiment and learning $w$ corresponding to the fixed prompt. In other words, $w$ is learned for calibration for every possible ICL prompt.
In Appendix Alg. 1, we introduce the recalibration algorithm employing temperature scaling. Additionally, we utilize the scaling-binning calibrator (Kumar et al., 2019), which fits a calibration function $w\in\mathcal{W}$ to the recalibration dataset: $\arg\min_{w}\sum_{(x_{i},y_{i})}\ell(w\cdot\mathcal{P}_{\theta}(x_{i}),y_{i})$, where $\ell$ is log-loss. Subsequently, the input space is partitioned into bins, ensuring an equal number of inputs in each bin (defaulting to 10 bins). Within each bin, the average of the $w$ values is computed and outputted for recalibration.
@@
Upon examination of Table 3 and Table 4, it is evident that none of the aforementioned strategies utilizing temperature scaling achieves satisfactory calibration performance. This finding contrasts with the well-established success of scaling confidence scores in the supervised learning setting, where it effectively reduces calibration errors (Guo et al., 2017). The fact that applying a post-processing calibration method, such as temperature scaling, cannot directly resolve the miscalibration issue suggests that ICL might have different properties compared to predictions from classical supervised learning models. On the other hand, the scaling-binning method demonstrates superior performance in our experiments, which successfully reduces calibration errors below $0.1$.
(a) Classification accuracy
(b) Calibration error
@@
Figure 3: Accuracy and calibration errors of base models LLaMA and Mistral, as well as fine-tuned variants. Reported Acc and ECE results are averaged across experiments conducted with $\{0,1,2,4,8\}$ shots.
Dataset Strategy 0-shot 1-shot 2-shot 3-shot 4-shot 8-shot Avg
SST-2 None 0.043 0.223 0.119 0.101 0.060 0.049 0.099
00-shot 0.043 0.216 0.082 0.074 0.047 0.057 0.087
$k$-shot 0.034 0.197 0.101 0.079 0.041 0.038 0.139
Fix $w$ 0.035 0.176 0.086 0.073 0.047 0.043 0.077
CB None 0.125 0.316 0.177 0.202 0.221 0.210 0.203
00-shot 0.015 0.252 0.162 0.217 0.217 0.199 0.209
$k$-shot 0.015 0.357 0.187 0.188 0.212 0.216 0.214
Fix $w$ 0.015 0.217 0.159 0.173 0.182 0.210 0.190
RTE None 0.108 0.110 0.142 0.122 0.128 0.120 0.122
00-shot 0.107 0.112 0.143 0.114 0.125 0.116 0.119
$k$-shot 0.108 0.115 0.136 0.112 0.126 0.125 0.120
Fix $w$ 0.101 0.082 0.097 0.068 0.076 0.097 0.088
AGNews None 0.089 0.057 0.071 0.121 0.085 0.123 0.090
00-shot 0.067 0.087 0.098 0.160 0.107 0.130 0.114
$k$-shot 0.083 0.074 0.059 0.109 0.073 0.082 0.079
Fix $w$ 0.080 0.074 0.080 0.091 0.073 0.080 0.080
@@
Table 3: ECE for different calibration strategies using temperature scaling (Guo et al., 2017) of base models LLaMA-2-7B across various shot settings.
Dataset Strategy 0-shot 1-shot 2-shot 3-shot 4-shot 8-shot Avg
SST-2 None 0.043 0.223 0.119 0.101 0.060 0.049 0.099
00-shot 0.015 0.062 0.055 0.060 0.062 0.057 0.052
$k$-shot 0.022 0.007 0.008 0.013 0.004 0.008 0.010
Fix $w$ 0.021 0.004 0.008 0.010 0.005 0.009 0.010
CB None 0.125 0.316 0.177 0.202 0.221 0.210 0.203
00-shot 0.122 0.130 0.121 0.086 0.083 0.119 0.110
$k$-shot 0.119 0.109 0.100 0.109 0.101 0.049 0.094
Fix $w$ 0.119 0.088 0.085 0.110 0.121 0.069 0.099
RTE None 0.108 0.110 0.142 0.122 0.128 0.120 0.122
00-shot 0.078 0.083 0.090 0.100 0.102 0.115 0.093
$k$-shot 0.089 0.084 0.089 0.095 0.101 0.112 0.096
Fix $w$ 0.077 0.086 0.092 0.100 0.108 0.117 0.099
AGNews None 0.089 0.057 0.071 0.121 0.085 0.123 0.090
00-shot 0.007 0.013 0.011 0.014 0.013 0.014 0.013
$k$-shot 0.001 0.009 0.015 0.018 0.005 0.005 0.009
Fix $w$ 0.015 0.019 0.019 0.005 0.008 0.017 0.013
Table 4: ECE for different calibration strategies using scaling-binning (Kumar et al., 2019) calibrator of base models LLaMA-2-7B across various shot settings.
(a) 00-shot
(b) $1$-shot
(c) $4$-shot
(d) $8$-shot
@@
Figure 4: Illustration of confidence distribution. The number of samples whose confidence is greater than a threshold on Commonsense QA.
(a) 00-shot
(b) $1$-shot
(c) $4$-shot
(d) $8$-shot
@@
The effect of fine-tuning. We show that vicuna, alpaca, and LLaMA2-Chat are all more accurate but less calibrated than their LLaMA counterpart backbones (Fig. 3), the margin is especially large for reasoning tasks and vicuna. Our finding indicates that fine-tuning might significantly degrade calibration, corroborating the evidence reported in GPT-4 (OpenAI, 2023), albeit it can greatly improve the reasoning accuracy. Our results provide evidence that though fine-tuning on carefully curated datasets can greatly improve question-answering performance, especially for hard tasks like reasoning problems, attention may need to be paid when assessing the calibration of those models’ predictions. Moreover, we include results of Mistral-7B (Jiang et al., 2023), a sparse Mixture of Experts (MoEs) architecture with sliding window attention. As a base model, it shows similar performance and calibration compared with LLaMA2-7B, indicating that our conclusion still holds for the model pre-trained with significantly different data and architecture. Comprehensive results and variance across different configurations are elaborated in Appendix Table 11.
@@
The findings, as detailed in Tab. 2, reveal certain insights: firstly, integrating labels within prompts significantly decreases uncertainty and enhances learning performance. The reason may be that it aids the model in understanding the label space, which leads to better classification outcomes. In contrast, simply repeating the context without labels does not lead to better outcomes. Secondly, the diversity of ICL examples in the prompt greatly affects performance, a potential explanation is it promotes better task learning (Pan, 2023). Those observations corroborate that ICL is performant when the number of ICL examples is large and they demonstrate consistent task properties. Importantly, the trade-off persists for different controlled scenarios, i.e. as we increase the number of ICL examples, models tend to be first more miscalibrated and then calibrated.
@@
Table 5: Norm of representation, entropy, and confidence of LLaMA-30B across three text classification datasets.